OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (callisto):

First, expand the terms... 3 − 5(x + 2) = 3(x + 1) + 6 3 - 5(x) -5 (2) = 3(x) + 1(3) +6 3 - 5x - 10 = 3x + 3 +6 Then, add the like terms (3-10) -5x = 3x + (3+6) -7 -5x = 3x +9 Then, move the term with variable x on one side and the constant term on the other side -7 - 5x +( 5x -9) = 3x + 9 +( 5x -9) -16 = 8x Divide both sides by the coefficient of x (which is 8 in this case) -16/8 = 8x/8 -2 = x Make x on the left ._. x=-2 Obviously, I'm not a good 'teacher'...

OpenStudy (callisto):

welcome~